OpenStudy (anonymous):

Can someone translate the percent word problem in a equation: If 5 is increased to 8, the increase is what percent of the original number

OpenStudy (jamesj):

If you increased 5 to 10, then I would have doubled the number 5 and this is a 100% increases. If you increased 5 to 7.5, then this is 50% increase, because 50% of 5 is 2.5 and 5 + 2.5 = 7.5 If you increased 5 to 6, then this is a 20% increase, because 5 + 1 = 6 and 1 is 20% of 5 If you increased 5 to 7, then this is a 40% increase, because 5 increased by 2, 5 + 2 = 7, and 2/5 = 40% If you increased 5 to x, then 5 was increased by \[\frac{x}{5} \times 100\%\] For example 5 to 9 is a \[\frac{4}{5} \times 100\% = 80 \% \ \text{increase}\] So here 5 increases by 8, 5 + 3 = 8, thus the percentage increase is \[\frac{3}{5} \times 100\% = ... what? \]

OpenStudy (anonymous):

Alright,then, thank you JameJ.

OpenStudy (jamesj):

So, what answer do you get?

OpenStudy (anonymous):

0.006

OpenStudy (jamesj):

No, that can't be right, it's far too small. But the 6 part is right.

OpenStudy (jamesj):

because we already know that 5 --> 7.5 is a 50% increase so 5 --> 8 has to be more than a 50% increase. We also know that 5 --> 9 is a 80% increase, so 5 --> 8 has to be less than a 80% increase.

OpenStudy (anonymous):

My bad ,I calculatored wrong by 5(100) so the actually number number will be 60 since, I I made a fraction of 3/5 with 100/ and multiply both to get 300/5 and reduce it to get 60.

OpenStudy (jamesj):

Yes, 60. Great.

OpenStudy (anonymous):

Thank, also I have another question: There were 1200 single women who participated in the poll. How many stated they would marry someone other than the perfect mate? on this pie bar graph: 34% said yes, 61% said no, and 5% said not sure. What is the first step of finding this answer?

OpenStudy (anonymous):

Well,it alright then, thank for your help today, Have a nice day.

OpenStudy (jamesj):

Oh, I didn't see this new question until now. Post it again on the left and someone, including maybe me, will help you.